On June 30, 2017, Wisconsin, Inc., issued $109,300 in debt and 22,800 new shares of its $10 par value stock to Badger Company owners in exchange for all of the outstanding shares of that company. Wisconsin shares had a fair value of $40 per share. Prior to the combination, the financial statements for Wisconsin and Badger for the six-month period ending June 30, 2017, were as follows Wisconsin Badger Revenues $ (1,001,000) $ (428,0ee) Expenses 741,000e 297,000 $(131,000) $ (277,000) Net income $ (260,000) Retained earnings, 1/1 Net income Dividends declared $ (815,000) (260,000) 103,508 $ (971,500) (131,000) Retained earnings, 6/30 Cash Receivables and inventory Patented technology (net) Equipment (net) $ (408,000) $ 122,000 172,000 321,000 637,000 $62,500 463,000 927,000 728,000 $ 2,180,500 $1,252,000 Total assets Liabilities Common stock Additional paid-in capital Retained earnings $ (579,000) (360,000) (270,000) $ (374,000) (200,000) (270,000) 408,000 971,500e Total liabilities and equities $ (2,180,500) $ (1,252,000)Wisconsin also paid $31,400 to a broker for arranging the transaction. In addition, Wisconsin paid $47,400 in stock issuance costs. Badgers equipment was actually worth $757,000, but its patented technology was valued at only $293,300. What are the consolidated balances for the following accounts? (Input all amounts as positive values) Accounts Amounts a. Net income b. Retained earnings, 1/1/17 c. Patented technology d. Goodwill e. Liabilities f. Common stock. g. Additional paid-in capital.
Answer #1

The shares that has been issued by Wisconsin will be recorded at fair value. Value of common stock = 16300 shares*$10 = 163,000. Additional paid in capital = 16300*(fair value - par value) = 16300*(40-10) = 489,000. Liabilities = 339900. Journal entry will be:

Dr Cr
Investment in Badger (value of debt and shares issued) 991,900
Common Stock (par value) 163,000
Additional Paid In Capital (excess over par value) 489,000
Liabilities 339,900

$34,700 was paid to the broker. This will be recorded as an expense.

Dr Cr
Professional Services Expense 34,700
Additional Paid In Capital 47,500
Cash 82,200

Calculation of allocation of acquisition date excess fair value:

Consideration transferred (fair value) for Badger Stock   991,900
less: Book Value of Badger 811,000
Fair Value in Excess of Book Value 180,900
Less: Excess fair value (undervalued equipment) 156,000
Add: Excess fair value (overvalued patented technology) 29,500
Goodwill 54,400

Calculations:

Book value of Badger = 1262000-451000 (total liabilities and equitues - total liabilities)

Undervalued equipment = 803000 - 647000 (actual worth - value as per balance sheet). Overvalued patented technology has been calculated similarly.

Consolidated balances:

1. Net Income: net income of Wisconsin - expense of profession service = 329,000 - 34,700 = $294,300

2. Retained earnings: 1,089,250

3. Patented technology: 951,000+274,500 = $1,225,500

4. Goodwill: 54,400 (calculations shown in the table above)

5. Liabilities: 586,000+339,900 (liabilities as calculated in Badger)+451,000 = $1,376,900

6. Common stock: 360,000+163,000 (investment in Badger) = $523,000

7. Additional paid in capital: 270,000+ 489,000 (Investment in Badger)-47500 (stock issuance costs) = $711,500